# Measurement-based Fast Quantum State Stabilization with Deep Reinforcement Learning

### v1

@@ 4.2 Reward
where $\overline{D}$ and $\underline{D}$ as the upper and lower bounds of the distance $D_{\rho_{t}}$, respectively, and $\overline{R}$ and $\underline{R}$ as the upper and lower bounds of the reward, respectively. In this way, we can set the numerical values of these four bounds to assign different rewards to different $D_{\rho_{t}}$. Additionally, we use $\mathfrak{e}$ and $\mathfrak{f}$ to regulate the slope of the reward curve. In this paper, we set $\mathfrak{e}=2$ and $\mathfrak{f}=10$. The different slopes imply that different distances have different importance for the DRL agent.
When the distance $D_{\rho_{t}}\leq 0.001$, the state is considered nearly ready for success, and the system deserves a positive reward. As shown in Figure 2(a), when the distance is between $0.001$ and 00 (i.e., $\overline{D}=0.001$ and $\underline{D}=0$), we set the reward boundaries to range from $1$ to $100$ (i.e., $\overline{R}=100$ and $\underline{R}=1$). The closer the distance is to 00, the larger the positive reward is given. The steepening slope of the reward curve as the system approaches the goal underscores the urgency of achieving the target state rapidly. This design enforces the agent’s commitment to swift and accurate goal attainment.
However, providing positive rewards only when the target is achieved can lead to the reward sparse problem in reinforcement learning [37]. In the absence of supplementary guidance to direct the agent towards the correct trajectory prior to engaging with the objective, the agent is likely to encounter substantial difficulties in accomplishing the mission solely through random exploration. To address this issue, we introduce a negative reward (penalty) at each step when the system is not in the vicinity of the target state. The punishment range is set to $[-0.1,0]$ as depicted in Figure 2(b). When $D_{\rho_{t}}=1$, a punishment of $-0.1$ is given. As the distance $D_{\rho_{t}}$ decreases, indicating that the system state is getting closer to the goal, the punishment decreases faster (represented by a steeper slope). When the distance reaches $0.001$, the previously described positive reward in Figure 2(a) takes over.
Remark 1: When the distance $D_{\rho_{t}}$ approaches $0.001$, we choose to gradually reduce the penalties rather than increase positive rewards. This decision is made to stabilize the DRL agent during training and mitigate the issues related to reward hacking [38]. For instance, if we were to provide progressively larger rewards as $D$ approaches $0.001$, the agent might be inclined to persist near $0.001$ until reaching the maximum number of steps in the episode, as this would yield simpler and higher cumulative rewards. By applying small penalties at each step, we encourage the agent to explore the state space more extensively. Our aim is to prevent the agent from becoming overly fixated on known high-reward regions while neglecting other unknown regions that may offer higher rewards. Instead, we prompt the agent to reach the goal promptly and iteratively improve its policy to enhance performance.
Additionally, recognizing the desirability of swiftly stabilizing the system to the target state, we introduce a small additional penalty associated with the current number of evolutionary steps. Prior to system stabilization at the target, an additional penalty proportional to the number of steps is applied. For instance, the first step incurs a penalty of $-1\times 10^{-6}$, the second step incurs $-2\times 10^{-6}$, and so forth.
In summation, the outlined approach encompasses the design of instantaneous rewards for each step. This reward design encourages the agent to explore and discover more possibilities while guiding it to reach the target state quickly, and effectively. The balance between positive rewards when close to the goal and negative rewards when away from it helps the agent to learn from both “good” and “bad” actions, leading to a more robust and effective learning process. We show the superiority of our design by simulating different reward functions in the Appendix.
@@
Figure 2: Reward Function. (a) When the $D_{\rho_{t}}$ is between $0.001$ and 00, we consider that the current system is almost approaching the target state. A smaller $D_{\rho_{t}}$ value corresponds to a larger positive reward. When $D_{\rho_{t}}=0$, the maximum reward $r=100$ is obtained. (b) When the $D_{\rho_{t}}$ is between $1$ and $0.001$, we consider that there is still distance between the current state and the target state. At $D_{\rho_{t}}=1$, the maximum negative reward (penalty) is imposed, and as the $D_{\rho_{t}}$ value decreases, the penalty decreases accordingly.
@@
$A^{\theta}(\rho_{t},a_{t})=Q(\rho_{t},a_{t})-V^{\phi}(\rho_{t})$ is known as the advantage function in the field of reinforcement learning, which is utilized to assess the desirability of taking a specific action $a_{t}$ at state $\rho_{t}$. $Q(\rho_{t},a_{t})=\sum_{t^{\prime}=t}^{m}\gamma^{t^{\prime}-t}r_{t^{\prime}}$ represents the action-value function, indicating the expected discounted reward for choosing action $a_{t}$ in state $\rho_{t}$, i.e., the cumulative sum of rewards until the end of the episode after executing this action. $r_{t^{\prime}}$ is the reward function in Eq. (7). The value of $\gamma$ lies between 00 and $1$, determining the emphasis on long-term rewards (close to $1$) or short-term rewards (close to 00). It effectively introduces a discounting mechanism for future rewards, thereby shaping the agent’s preference for future reward consideration when making decisions. $V^{\phi}(\rho_{t})$ is referred to as the state-value function (or baseline) and is modeled by a neural network with the same structure as the policy network but with different parameters $\phi$. It is primarily employed to approximate the discounted rewards from state $\rho_{t}$ to the end of an episode. Specifically, if the current state is $\rho_{t}$, and for all possible actions $a_{t}^{(1)},a_{t}^{(2)},\dots,a_{t}^{(\dots)}$, they correspond to discounted rewards $Q(\rho_{t},a_{t}^{(1)}),Q(\rho_{t},a_{t}^{(2)}),\dots,Q(\rho_{t},a_{t}^{(\dots)})$. As $V^{\phi}(\rho_{t})$ represents the expected value of the discounted rewards at $\rho_{t}$, we can use $Q(\rho_{t},a_{t}^{(1)}),Q(\rho_{t},a_{t}^{(2)}),\dots,Q(\rho_{t},a_{t}^{(\dots)})$ as features to approximate the value of $V^{\phi}(\rho_{t})$, representing the expected value of rewards in state $\rho_{t}$. When $A^{\theta}(\rho_{t},a_{t})>0$, action $a_{t}$ is considered better than average and is worth increasing the probability of being chosen in subsequent iterations while decreasing the probability otherwise.
4.3 Core Algorithmic Ideas for DRL Agents
To maximize $J(\theta)$, which indicates that our chosen policy parameters $\theta$ can lead to higher average rewards, we adopt the well-known gradient descent method. Thus, we take the derivative of the expected reward $J(\theta)$ in Eq. (10), resulting in the expression shown in
$\nabla J(\theta)$ $=\sum_{\tau}R(\tau)\nabla p_{\theta}(\tau)$
$=\sum_{\tau}R(\tau)p_{\theta}(\tau)\nabla\log\,p_{\theta}(\tau)$
We use $\nabla f(x)=f(x)\nabla\log f(x)$ to derive the second row of Eq. (11). The last approximate equation is a result of practical gradient computations, where instead of calculating the expected reward for an entire trajectory, rewards contributed by each individual state-action pair ($\rho,a$) are computed separately. These individual rewards are then summed up to obtain the total cumulative reward for the optimization process. The direction of the policy update $\pi_{\theta}(a|\rho)$ is biased towards favoring state-action pairs that contribute to higher cumulative rewards within the sequence. For instance, if an action $a$ executed in state $\rho$ leads to a positive cumulative discounted reward, the subsequent update will enhance the probability of choosing action $a$ in state $\rho$, while diminishing the likelihood of selecting other actions. The update equation for the parameters $\theta$ is as follows:
$\theta=\theta+\eta\nabla J(\theta),$ (12)
where $\eta$ is the learning rate.
Once the policy $\pi_{\theta}(a|\rho)$ is updated, it necessitates the reacquisition of training data prior to the subsequent policy update. This arises due to the alteration in the probability distribution $p_{\theta}(\tau)$ brought about by the modified policy. Following data sampling, the parameter $\theta$ undergoes refinement, leading to the discarding of all prior data. Subsequent parameter updates mandate the collection of fresh data, constituting the fundamental principle underlying the conventional PG algorithm. However, in the context of quantum systems, the process of sampling system information is often characterized by time-intensive and computationally demanding operations. For instance, after each measurement, a classical computer is requisitioned to solve the SME (1) to ascertain the system’s state in the subsequent moment. This inability to reutilize previously acquired data contributes to a protracted training process. To address this challenge, an additional strategy $\pi_{\theta^{\prime}}$ is introduced, mirroring the architecture of $\pi_{\theta}(a|\rho)$. Instead of directly engaging with the environment for data gathering, the primary agent $\pi_{\theta}$ employs the auxiliary agent $\pi_{\theta^{\prime}}$ to interact with the environment and accumulate data. The objective is to subsequently utilize this data to train $\pi_{\theta}$ multiple times, effectively reducing the computational and resource demands for data collection. Ensuring the consistency of data sampled by $\pi_{\theta^{\prime}}$ with that of $\pi_{\theta}$, importance sampling [39] is introduced to facilitate this synchronization process. This approach contributes to enhancing data reuse and the overall efficiency of the training procedure. Eq. (11) is updated as:
$\nabla J(\theta)$
Here, all the state-action pairs $(\rho_{t},a_{t})$ (or alternatively, all trajectories $\tau\in\mathbb{D}$) are sampled from $\pi_{\theta^{\prime}}$, where ${p_{\theta}(\rho_{t},a_{t})}/{p_{\theta^{\prime}}(\rho_{t},a_{t})}$ is the importance weight, dynamically adjusting the weight of the data sampled by $\pi_{\theta^{\prime}}$ in real-time to more accurately estimate the expected value under the target policy $\pi_{\theta}$.
The corresponding objective function from Eq. (13) can be calculated as:
$J_{\rm PPO}^{\theta^{\prime}}(\theta)$
$J_{\rm PPO}^{\theta^{\prime}}(\theta)$
  • One set of main policy parameters $\theta$, which is updated every time.
4.4 Training
@@
Neural Network Architecture: Each policy $\pi_{\theta}$ is represented by a neural network that maps a given state $\rho$ to a probability distribution over actions $a$. The action distribution is modeled as a gaussian distribution. The input layer is processed by two fully connected hidden layers, each with $128$ neurons, accompanied by a linear output layer with the same dimension as the action space ($2$ in this paper). All hidden layers use the Tanh activation function. The value function $V^{\phi}(\rho_{t})$ is composed of a similar neural network architecture, with the only difference being that the output layer is a single linear unit used to estimate the state-value function. The value function $V^{\phi}(\rho_{t})$ is estimated using the temporal difference (TD) method [42]. Then, the generalized advantage estimator (GAE) [43] is employed to compute the advantage function in Eq. (8), which is subsequently used in Eq. (15) to calculate the gradient for updating the policy $\pi_{\theta}$.
@@
  • When the distance $D_{\rho_{t}}\in[0,0.001]$ for $10$ consecutive measurements, multiple measurements verifying that the system is in the target state are considered mission complete.
  • In a specific system, the maximum training time for a trajectory is set to a fixed value. For example, for the two-qubit state stabilization problem in section 5.1, we set the maximum training time $T=20$ arbitrary units (a.u.). When the evolution time reaches $20$ a.u., regardless of whether it has converged to the goal or not, the training trajectory is halted. This approach not only greatly saves training time but also significantly reduces the issue of overfitting.
@@
           compute advantage $A^{\theta}$ in Eq. (8) using current $V^{\phi}$ and GAE
@@
        Calculate the gradient to update the policy parameters $\theta$ via Eq. (16)
@@
5 Numerical Simulation
5.1 Two-Qubit System
We first consider a system of two qubits in a symmetric dispersive interaction with an optical probe as proposed in [45]. We consider using the DRL control scheme mentioned above to address the two-qubit entangled state preparation problem for arbitrary initial quantum states. Denote the Pauli matrices
@@ 5.1 Two-Qubit System
Control Hamiltonians $H_{1}$ and $H_{2}$ are
$H_{1}=\sigma_{y}\otimes I_{2}\\ ={\begin{bmatrix}0&0&-i&0\\ 0&0&0&-i\\ i&0&0&0\\ 0&i&0&0\end{bmatrix}},$
$H_{2}=I_{2}\otimes\sigma_{y}={\begin{bmatrix}0&-i&0&0\\ i&0&0&0\\ 0&0&0&-i\\ 0&0&i&0\end{bmatrix}},$
respectively. Angular momentum operator
$c=\sigma_{z}\otimes I_{2}+I_{2}\otimes\sigma_{z}\\ ={\begin{bmatrix}2&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&-2\end{bmatrix}}.$
@@
$\rho_{d}={\begin{bmatrix}0&0&0&0\\ 0&0.5&0.5&0\\ 0&0.5&0.5&0\\ 0&0&0&0\end{bmatrix}},$
which is a symmetric two-qubit state.
@@
We utilize the previously summarized PPO algorithm to train the DRL agent. For the training trajectories, we set a time interval of $\Delta t=0.001$ a.u. for each measurement step, with a maximum evolution time of $T=20$ a.u., corresponding to a maximum of $20,000$ steps. At each step, the DRL agent interacts with the environment, obtaining system information to generate control signals, which are then stored for iterative updates of the policy. The total number of training steps is set to $10^{7}$. On our computer, training takes approximately $50$ minutes. However, as our primary focus is on the performance of the trained agent, the training duration is not of critical importance. In practical experiments, it is feasible to pre-train the agent; once trained, the agent can be directly utilized.
@@
In order to evaluate its performance, we test the proposed strategy on $50$ randomly selected distinct initial quantum states $\rho_{0}$ (corresponding to different initial distances $D_{\rho_{0}}$, as indicated by the blue line in Figure 3). The light blue lines represent the evolution trajectory of a specific initial state with respect to the target state, averaged over $50$ different trajectories under varying environmental noise, while the dark blue line depicts the average evolutionary trajectory of all the different initial states. A distance of $D_{\rho_{t}}=0$ indicates the system stabilizing at the target state. It is worth noting that the well-trained agent successfully stabilizes any initial state to the target state. Furthermore, for comparison with the Lyapunov method mentioned in [45], we retain the same $50$ sets of randomly selected initial states and obtain the orange trajectories in Figure 3 using the Lyapunov method. It can be observed that the control signals generated by the DRL agent outperform the Lyapunov method. Assuming we take the time when the distance $D_{\rho_{t}}$ becomes less than $0.001$ as the required evolution time of the system, under the guidance of the DRL agent, the average evolution time is $5.64$ a.u., while the average time using the Lyapunov method is $14.36$ a.u.. The DRL’s average stabilization time is improved by $60.72\%$ over the Lyapunov method. This indicates that our DRL approach successfully stabilizes these quantum states to the target state faster than the Lyapunov method.
@@
Figure 3: Evolution of the distance $D_{\rho_{t}}$ for $50$ random initial states stabilized to the target Two-Qubit state under the control of the DRL agent (blue) and Lyapunov method (orange). The average stabilization time to the target under DRL control is $5.64$ a.u., while the Lyapunov method requires an average time of $14.36$ a.u.. (The stabilization time is defined as the time when the distance $D_{\rho_{t}}\leq 0.001$.) The light blue (orange) lines represent the average evolution trajectories for different initial states, and the dark blue (orange) line represents the average trajectory across all different initial states.
5.2 GHZ State

### v2

@@ IV-B Reward
where $\overline{D}$ and $\underline{D}$ are the upper and lower bounds of the distance $D_{\rho_{t}}$, $\overline{R}$ and $\underline{R}$ are the upper and lower bounds of the reward, and $\mathfrak{e}$ and $\mathfrak{f}$ are parameters that regulate the slope of the reward curve.
In general, this reward function is motivated from the inverse proportional function (the relation between $r_{t}$ and $D_{\rho_{t}}$ when other parameters in (7) are fixed), which ensures that the reward value increases when the distance $D_{\rho_{t}}$ decreases. More specifically, the bounds $\overline{D}$ and $\underline{D}$ are designed to ensure $D_{\rho_{t}}\in[\underline{D},\overline{D}]$. The distance is mapped to a reward value constrained within the range $[\underline{R},\overline{R}]$. The two tunable parameters, $\mathfrak{f}$ and $\mathfrak{e}$ are used to adjust the steepness when $D_{\rho_{t}}$ is approaching $\underline{D}$ and $\overline{D}$, respectively.
The detailed effects of different $\mathfrak{e}$ and $\mathfrak{f}$ are plotted in Fig. 2. For the given bounds $\underline{D},\overline{D},\underline{R}$, and $\overline{R}$, when $\mathfrak{e}<\mathfrak{f}$, the reward curve near $\underline{D}$ is steeper than that near $\overline{D}$. This indicates that the rate of increase in the reward value is more pronounced as $D_{\rho_{t}}$ approaches $\underline{D}$. Conversely, when $\mathfrak{e}>\mathfrak{f}$, the trend is reversed.
Figure 2: Example Reward Curves for Different Parameter Combinations of $\mathfrak{e}$ and $\mathfrak{f}$.
In most existing work, the upper and lower bounds of $D_{\rho_{t}}$ are fixed at $\overline{D}=1$ and $\underline{D}=0$, respectively. The reward function in Eq. (7), however, offers the flexibility to adjust these bounds, as well as the corresponding reward values $\overline{R}$ and $\underline{R}$. This flexibility enables the division of the state space into multiple regions, which is particularly useful for addressing complex problems. Moreover, by adjusting $\overline{D}$ and $\underline{D}$, it becomes possible to assign positive reward values in certain regions of the state space and negative reward values (penalties) in others. This feature allows for more efficient and effective optimization, resulting in improved performance.
In this paper, we divide the state space into two regions using a partition parameter $d$. These two regions are denoted as Proximity Zone ($D_{\rho_{t}}<d$) and Exploration Zone ($D_{\rho_{t}}\geq d$). The same reward function in the formula (7) will be used in both of these two zones, and the corresponding bounds, $\underline{D},\overline{D}$ are determined by the partition parameter $d$. This division provides more flexibility to balance the reward and the penalty in the whole state space. One possible curve of the reward function in these two zones are shown in Fig.  3.
@@
Figure 3: Reward Function. (a) When the $D_{\rho_{t}}$ is between 00 and $d$, we consider that the current system is almost approaching the target state. A smaller $D_{\rho_{t}}$ value corresponds to a larger positive reward. When $D_{\rho_{t}}=0$, the maximum reward $r=\overline{R}_{P}$ is obtained. (b) When the $D_{\rho_{t}}$ is between $d$ and $1$, we consider that there is still distance between the current state and the target state. At $D_{\rho_{t}}=1$, the maximum negative reward (penalty) is imposed, and as the $D_{\rho_{t}}$ value decreases, the penalty decreases accordingly.
In the Proximity Zone ($D_{\rho_{t}}<d$), the state is considered close to success. Positive rewards encourage the agent to converge rapidly to the target state. The distance bounds in this region are $\overline{D}=\overline{D}_{P}=d$ and $\underline{D}=\underline{D}_{P}=0$, while the reward bounds are $\overline{R}_{P}>\underline{R}_{P}\geq 0$. As shown in Fig. 3(a), the reward function ensures that the slope of the curve increases as $D_{\rho_{t}}$ approaches zero, emphasizing rapid stabilization near the target state.
In the Exploration Zone ($D_{\rho_{t}}\geq d$), the system is considered far from the target state, and penalties are applied to guide exploration. The boundaries of this zone are $\overline{D}=\overline{D}_{E}=1$ and $\underline{D}=\underline{D}_{E}=d$, with the reward range satisfying $\underline{R}_{E}<\overline{R}_{E}\leq 0$. As illustrated in Fig. 3(b), the penalty decreases as $D_{\rho_{t}}$ approaches $d$, with the slope of the curve ensuring a progressively faster reduction. This mechanism encourages the state to transition into the Proximity Zone.
Remark 1: In the Exploration Zone, when the distance $D_{\rho_{t}}$ approaches $d$, we reduce penalties rather than introduce large positive rewards. This approach avoids potential reward hacking [33], where the agent might exploit increasing rewards by staying near $D_{\rho_{t}}\approx d$ without fully reaching the target. By applying small penalties at each step, the agent is encouraged to explore more broadly while progressively improving its policy.
Additionally, a small penalty proportional to the number of steps taken before stabilization is introduced to encourage efficiency. For example, the first step incurs a penalty of $-1\times 10^{-6}$, the second step $-2\times 10^{-6}$, and so forth. This step-based penalty discourages unnecessary delays and motivates the agent to achieve stabilization promptly.
The proposed reward function adjusts based on the distance $D_{\rho_{t}}$, with steeper slopes in both regions as the system approaches the target state. This design encourages exploration in the Exploration Zone while driving efficient convergence in the Proximity Zone. We show the superiority of our design by simulating different reward functions in Section VI.
@@
$A^{\theta}(\rho_{t},a_{t})=Q(\rho_{t},a_{t})-V^{\phi}(\rho_{t})$ is known as the advantage function in the field of RL, which is utilized to assess the desirability of taking a specific action $a_{t}$ at state $\rho_{t}$. $Q(\rho_{t},a_{t})=\sum_{t^{\prime}=t}^{m}\gamma^{t^{\prime}-t}r_{t^{\prime}}$ represents the action-value function, indicating the expected discounted reward for choosing action $a_{t}$ in state $\rho_{t}$, i.e., the cumulative sum of rewards until the end of the episode after executing this action. $r_{t^{\prime}}$ is the reward function in (7). The value of $\gamma$ lies between 00 and $1$, determining the emphasis on long-term rewards (close to $1$) or short-term rewards (close to 00). It effectively introduces a discounting mechanism for future rewards, thereby shaping the agent’s preference for future reward consideration when making decisions. $V^{\phi}(\rho_{t})$ is referred to as the state-value function (or baseline) and is modeled by a neural network with the same structure as the policy network but with different parameters $\phi$. It is primarily employed to approximate the discounted rewards from state $\rho_{t}$ to the end of an episode. Specifically, if the current state is $\rho_{t}$, and for all possible actions $a_{t}^{(1)},a_{t}^{(2)},\dots,a_{t}^{(\dots)}$, they correspond to discounted rewards $Q(\rho_{t},a_{t}^{(1)}),Q(\rho_{t},a_{t}^{(2)}),\dots,Q(\rho_{t},a_{t}^{(\dots)})$. As $V^{\phi}(\rho_{t})$ represents the expected value of the discounted rewards at $\rho_{t}$, we can use $Q(\rho_{t},a_{t}^{(1)}),Q(\rho_{t},a_{t}^{(2)}),\dots,Q(\rho_{t},a_{t}^{(\dots)})$ as features to approximate the value of $V^{\phi}(\rho_{t})$, representing the expected value of rewards in state $\rho_{t}$. When $A^{\theta}(\rho_{t},a_{t})>0$, action $a_{t}$ is considered better than average and is worth increasing the probability of being chosen in subsequent iterations while decreasing the probability otherwise.
IV-C Training
The essence of employing DRL to address MDPs is centered on the training regimen of the agent, which is governed by the policy $\pi_{\theta}(a|\rho)$ This study implements the PPO algorithm, a model-free policy gradient method within the domain of DRL. PPO incorporates a “clip ratio” mechanism that limits the extent of policy updates, thereby enhancing the stability and reliability of the learning process. This approach mitigates the requirement for extensive environmental sampling, which is a common feature of traditional policy gradient techniques, and consequently, it improves the efficiency of the training phase. The PPO algorithm operates with three sets of network parameters: the primary policy parameters, policy parameter duplicates, and value network parameters. These parameters are instrumental in the iterative policy refinement and state value estimation processes. The comprehensive algorithmic details and the underlying mathematical formalism are presented in the Appendix.
@@ IV-C Training
Reward Function Parameters: We set $d=0.001$, which is the partition parameter of the reward function mentioned in Section IV-B. This means that when the distance $D_{\rho_{t}}$ is less than $0.001$, the system is considered to be close to the target and receives a positive reward. In this case, the fidelity of the system state can easily exceed $0.999$. Ideally, the smaller the value of $d$, the higher the control fidelity of the trained agent. However, as $d$ decreases, the training process becomes increasingly challenging and time consuming. Therefore, the choice of $d$ should strike a balance between achieving high fidelity and managing training complexity.
The upper and lower bounds of the reward in the Proximity Zone, $\overline{R}_{P}$ and $\underline{R}_{P}$ , are set to $100$ and $1$, respectively, encouraging the system to get as close as possible to the target. In the Exploration Zone, the upper and lower bounds of $\overline{R}_{E}$ and $\underline{R}_{E}$ are set to 00 and $-0.1$, respectively, maintaining a penalty without overly punishing the agent.
We set $\mathfrak{e}=2$ and $\mathfrak{f}=10$, emphasizing a steeper reward escalation as $D_{\rho_{t}}$ approaches $\underline{D}$.
@@
  • If the system is measured continuously for $10$ iterations, and the distance $D_{\rho_{t}}$ remains within the interval $[0,d]$ for all measurements, the system is considered to have reached the target state with high fidelity. At this point, the task is concluded.
  • In a specific system, the maximum training time for a trajectory is set to a fixed value. For example, for the two-qubit state stabilization problem in Section V-A, we set the maximum training time $T=20$ arbitrary units (a.u.). When the evolution time reaches $20$ a.u., regardless of whether it has converged to the goal or not, the training trajectory is halted. This approach not only greatly saves training time but also significantly reduces the issue of overfitting.
@@
           compute advantage $A^{\theta}$ in (8) using current $V^{\phi}$ and GAE
@@
        Calculate the gradient to update the policy parameters $\theta$ via (26)
@@
V Numerical Simulation
V-A Two-Qubit System
We consider a two-qubit system in a symmetric dispersive interaction with an optical probe, as described in [36]. The system’s dynamics are governed by the SME in (1), where we utilize a DRL control scheme to stabilize the system to a target entangled state from arbitrary initial states. Denote the Pauli matrices
@@ V-A Two-Qubit System
Control Hamiltonians in (1) are chosen as
$H_{1}=\sigma_{y}\otimes I_{2},\quad\text{and}\quad H_{2}=I_{2}\otimes\sigma_{y},$ (9)
which implies that two control channels are applied to this quantum system. And the forms of $H_{1}$ and $H_{2}$ represent $\sigma_{y}$ rotations on the first and second qubits, respectively, enabling independent control over each qubit.
The observable operator is chosen as:
$c=\sigma_{z}\otimes I_{2}+I_{2}\otimes\sigma_{z},$ (10)
which corresponds to the measurement of $\sigma_{z}$-like observable for the two qubits.
@@
$\rho_{d}={\begin{bmatrix}0&0&0&0\\ 0&0.5&0.5&0\\ 0&0.5&0.5&0\\ 0&0&0&0\end{bmatrix}},$ (11)
which is a symmetric two-qubit Bell state.
@@
We utilize the previously summarized PPO algorithm to train the DRL agent. For the training trajectories, we set a time interval of $\Delta t=0.001$ a.u. for each measurement step, with a maximum evolution time of $T=20$ a.u., corresponding to a maximum of $20,000$ steps. At each step, the DRL agent interacts with the environment, obtaining system information to generate control signals, which are then stored for iterative updates of the policy. The total number of training steps is set to $10^{7}$. On the computer with configuration in Section IV-C, the training process requires approximately $50$ minutes. However, the primary focus of this study is the performance of the trained agent rather than the optimization of training time. In practical implementations, employing a pre-trained agent is feasible. Once the target state is specified, the agent does not require repeated updates and can be directly utilized after training is completed.
@@
In order to evaluate its performance, we test the proposed strategy on $50$ randomly selected distinct initial quantum states $\rho_{0}$ (corresponding to different initial distances $D_{\rho_{0}}$, as indicated by the blue line in Fig. 4). The light blue lines represent the evolution trajectory of a specific initial state with respect to the target state, averaged over $50$ different trajectories under varying environmental noise, while the dark blue line depicts the average evolutionary trajectory of all the different initial states, i.e., the average trajectory of $2500$ different evolutionary trajectories. Smaller values of the distance $D_{\rho_{t}}$ indicate that the system is closer to the target state. It is worth noting that a trained agent can successfully stabilize any initial state to the target state with high fidelity. Furthermore, for comparison with the Lyapunov method mentioned in [36], we retain the same $50$ sets of randomly selected initial states and obtain the orange trajectories in Fig. 4 using the Lyapunov method. It can be observed that the control signals generated by the DRL agent outperform the Lyapunov method. Assuming that the time when the distance $D_{\rho_{t}}$ is less than $0.001$ is the evolution time of the system, the average evolution time of the $2500$ trajectories under the guidance of the DRL agent is $4.59$ a.u., while the average time using the Lyapunov method is $5.86$ a.u.. The DRL’s average stabilization time is improved by $22\%$ over the Lyapunov method. This indicates that our DRL approach successfully stabilizes these quantum states to the target state faster than the Lyapunov method.
@@
Figure 4: Evolution of the distance $D_{\rho_{t}}$ for $50$ random initial states stabilized to the target Two-Qubit state under the control of the DRL agent (blue) and Lyapunov method (orange). The average stabilization time to the target under DRL control is $4.59$ a.u., while the Lyapunov method requires an average time of $5.86$ a.u.. (The stabilization time is defined as the time when the distance $D_{\rho_{t}}\leq 0.001$.) The light blue (orange) lines represent the average evolution trajectories for different initial states, and the dark blue (orange) line represents the average trajectory across all different initial states.
V-B GHZ State